OpenStudy (anonymous):

determine whether the given number are solution of the inequalities x-3>= 9;-6,0,4,16. Is -6 a solution and is 0 a solution

OpenStudy (anonymous):

x-3 >= 9 x >= 12

OpenStudy (anonymous):

So the solution is any number bigger or equal to 12

OpenStudy (anonymous):

Plug in your number for x and see if it's true. \[x-3 \ge 9, x=-6 \rightarrow -6-3 \ge 9 ?\] -9 is not greater than or equal to 9 so -6 is not a possible solution.